OpenStudy (anonymous):

John has 15 more baseball cards than Michael. Together they have a total of 55 cards. How many cards does Michael have? Let M = the number of baseball cards Michael has.

OpenStudy (jackellyn):

M= J +15 M+J =55 Solve for J in the first equation m= j +15 m- 15 = j Substitute in second equation m + (m-15) = 55 m + m-15=55 2m - 15 = 55 2m = 70 m=35